# Bayesian Inference General Procedures for A Single-subject Test Study

### v2

@@ Acknowledgments
This work was supported by the High Performance Computing Cluster at University of Kent.
@@ Appendix A Appendix
Let $t(z|\nu)$ and $T(z|\nu)$ be the probability density function and cumulative distribution function of the Student’s $t$-distribution, respectively, for $\nu\geq 1$, then $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ can be approximated by $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$, where $\sigma_{\nu}=1.5536$ if $\nu>2400$. If $1\leq\nu\leq 2400$, then
$\sigma_{\nu}$ $=0.00000543(\log(\nu))^{7}-0.00016303(\log(\nu))^{6}+0.00199613(\log(\nu))^{5}-0.01285016(\log(\nu))^{4}$ (5)
$\quad+0.04631303(\log(\nu))^{3}-0.08761023(\log(\nu))^{2}+0.05036188\log(\nu)+1.62021189$
@@
given $\nu_{i}$, we use grid search method to minimise $KL(\sigma(\nu_{i}))$ and obtain the optimal scale $\hat{\sigma}(\nu_{i})$. Let $\sigma_{j,\nu_{i}}=1+0.0002j$, $j=0,1,\ldots,5000$, then $\hat{\sigma}_{\nu_{i}}=\operatorname*{arg\,min}_{\sigma_{j,\nu_{i}}}KL(\sigma_{j,\nu_{i}})$ where $KL(\sigma_{j,\nu_{i}})$ is evaluated using the Monte Carlo method. Finally, we carry out the polynomial fit with polynomial degree $p=2,3,\ldots,7$ based on the optimal scale $\hat{\sigma}(\nu_{i})$ and $\log(\nu_{i})$, $i=1,2,\ldots,245$Equation 5 is the best polynomial fit with polynomial degree $p=7$ based on Bayes information criterion, see Figure 6. Secondly, when $\nu>2400$, $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ can be approximated by $\phi(z)[\Phi(z)]^{-1/2}[1-\Phi(z)]^{-1/2}/\pi$. According to Bayes and Branco (2007), $\phi(z)[\Phi(z)]^{-1/2}[1-\Phi(z)]^{-1/2}/\pi$ can be further approximated by $N(0,\sigma^{2}_{\nu})$ where the optimal scale $\sigma_{\nu}$ chosen by total variational distance is 1.5536. Note that Bayes and Branco (2007) used the optimal scale $\sigma_{\nu}$ chosen by the Kullback-Leibler divergence (defined in Equation 7), the optimal scale $\sigma_{\nu}$ is 1.54 for normal distribution. However, our simulation result shows that the total variational distance criterion performs better than the Kullback-Leibler divergence criterion for Student’s $t$ with a small degree of freedom, see Figure 7. Therefore, we employ the criterion of total variational distance to choose the optimal scale $\sigma_{\nu}$ in $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$ to approximate $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ throughout this paper. Figure 8 shows the density $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ and its approximation $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$ with $\sigma_{\nu}$ in Equation 5 for $\nu=1,5,10,100$.
@@
Figure 7: Comparison of approximation based on Kullback-Leibler divergence and total variational distance when the degree of freedom $\nu=2$. The blue line represents the density of $f$, the green line represents the density of approximation $g$ based on total variational distance, the orange line is the density of approximation $g$ based on Kullback-Leibler divergence.
@@
Let $f(z|\alpha,\nu)$ be the non-linear skewed Student’s $t$ density function, defined as follows:
@@
By definition of Jeffery’s prior, we have $\boldsymbol{\pi}(\alpha|\nu)\propto\sqrt{I_{\alpha\alpha}}$, where $I_{\alpha\alpha}$ is the Fisher information of $\alpha$ in the non-linear skewed Student’s $t$ density function $f(z|\alpha,\nu)$. According to Branco et al. (2013, Proposition 2), the Fisher information of $\alpha$ in $f(z|\alpha,\nu)$ is given by
@@
$H_{6}$ $=\,_{2}F_{1}\left(\frac{1}{2},\nu+2;\frac{\nu+5}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right),$
which completes the proof. ∎
Goodness-of-fit for Skewed Student’s t Distribution
For simplicity, let $x_{1},x_{2},\ldots,x_{n}$ be the observations of $n$ control subjects. We denote $F(x|\alpha,\nu,\xi,\omega)$ as the cumulative function of this distribution where The parameters $\alpha,\nu,\xi,\omega$ are unknown. The definition of skewed Student’s $t$ distribution are defined in Section 2.2. To test the null hypothesis $H_{0}:x_{1},x_{2},\ldots,x_{n}$ are random samples from the skewed Student’s $t$ distribution, we propose the following testing process:
@@
More details on the testing procedures and critical points can be found in Chen and Balakrishnan (1995). If we accept the null hypothesis $H_{0}$ at some significance level, then we can proceed with the comparison procedures between a single-subject and the control group described in the next section. If $H_{0}$ is rejected, one can employ a non-parametric method to compare a single-subject with the control group. However, the discussion of the non-parametric method is beyond the scope of this paper, we plan to extend this approach in future work.
@@
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.0874 0.0788 0.0788 0.2214 0.0473
100 0.0677 0.0630 0.0630 0.1920 0.0389
200 0.0595 0.0571 0.0571 0.1972 0.0223
400 0.0533 0.0520 0.0520 0.1870 0.0144
TPR 50 0.9631 0.9574 0.9574 0.8090 0.9190
100 0.9820 0.9802 0.9802 0.8393 0.9571
200 0.9974 0.9968 0.9968 0.8586 0.9790
400 0.9983 0.9982 0.9982 0.8564 0.9819
ACC 50 0.9379 0.9393 0.9393 0.7938 0.9359
100 0.9572 0.9586 0.9586 0.8237 0.9591
200 0.9689 0.9698 0.9698 0.8307 0.9784
400 0.9725 0.9731 0.9731 0.8347 0.9837
Table 7: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under one-sided test (less). Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.0003 0.0002 0.0002 0.4511 0.0030
100 0.0000 0.0000 0.0000 0.4654 0.0044
200 0.0000 0.0000 0.0000 0.4699 0.0027
400 0.0000 0.0000 0.0000 0.4701 0.0021
TPR 50 0.6402 0.5907 0.5908 0.8229 0.8858
100 0.5445 0.5197 0.5198 0.7792 0.9206
200 0.5824 0.5693 0.5694 0.8491 0.9628
400 0.5389 0.5323 0.5323 0.8197 0.9670
ACC 50 0.8199 0.7953 0.7953 0.6859 0.9414
100 0.7722 0.7599 0.7599 0.6569 0.9581
200 0.7912 0.7847 0.7847 0.6896 0.9800
400 0.7694 0.7661 0.7661 0.6748 0.9825
Table 8: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under one-sided test(greater). Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
@@
Table 9: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under two-sided test. The underlying density is normal density function. Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
@@
Figure 10: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=1$ and $\nu=10$.  Figure 10(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.
@@
Figure 11: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=1$ and $\nu=5$.  Figure 11(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.
@@
Figure 12: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=2$ and $\nu=5$.  Figure 12(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.

### v4

@@ Acknowledgments
This KTP project (ID:10053865, KTP13481) received financial support from UKRI through Innovate UK and Innovision IP Ltd. This work was supported by the High-Performance Computing Cluster at the University of Kent.
@@ Appendix A Appendix
Let $t(z|\nu)$ and $T(z|\nu)$ be the probability density function and cumulative distribution function of the Student $t$-distribution, respectively, for $\nu\geq 1$, then $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ can be approximated by $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$, where $\sigma_{\nu}=1.5536$ if $\nu>2400$. If $1\leq\nu\leq 2400$, then
$\sigma_{\nu}$ $=0.00000543(\log(\nu))^{7}-0.00016303(\log(\nu))^{6}+0.00199613(\log(\nu))^{5}$ (5)
$\quad-0.01285016(\log(\nu))^{4}+0.04631303(\log(\nu))^{3}-0.08761023(\log(\nu))^{2}$
$\quad+0.05036188\log(\nu)+1.62021189$
@@
given $\nu_{i}$, we use grid search method to minimise $KL(\sigma(\nu_{i}))$ and obtain the optimal scale $\hat{\sigma}(\nu_{i})$. Let $\sigma_{j,\nu_{i}}=1+0.0002j$, $j=0,1,\ldots,5000$, then $\hat{\sigma}_{\nu_{i}}=\operatorname*{arg\,min}_{\sigma_{j,\nu_{i}}}KL(\sigma_{j,\nu_{i}})$ where $KL(\sigma_{j,\nu_{i}})$ is evaluated using the Monte Carlo method. Finally, we carry out the polynomial fit with polynomial degree $p=2,3,\ldots,7$ based on the optimal scale $\hat{\sigma}(\nu_{i})$ and $\log(\nu_{i})$, $i=1,2,\ldots,245$Equation 5 is the best polynomial fit with polynomial degree $p=7$ based on Bayes information criterion, see Figure 7. Secondly, when $\nu>2400$, $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ can be approximated by $\phi(z)[\Phi(z)]^{-1/2}[1-\Phi(z)]^{-1/2}/\pi$. According to Bayes and Branco (2007), $\phi(z)[\Phi(z)]^{-1/2}[1-\Phi(z)]^{-1/2}/\pi$ can be further approximated by $N(0,\sigma^{2}_{\nu})$ where the optimal scale $\sigma_{\nu}$ chosen by total variational distance is 1.5536. Note that Bayes and Branco (2007) used the optimal scale $\sigma_{\nu}$ chosen by the Kullback-Leibler divergence (defined in Equation 7), the optimal scale $\sigma_{\nu}$ is 1.54 for Normal distribution. However, our simulation result shows that the total variational distance criterion performs better than the Kullback-Leibler divergence criterion for Student $t$ with a small degree of freedom, see Figure 8. Therefore, we employ the criterion of total variational distance to choose the optimal scale $\sigma_{\nu}$ in $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$ to approximate $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ throughout this paper. Figure 9 shows the density $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ and its approximation $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$ with $\sigma_{\nu}$ in Equation 5 for $\nu=1,5,10,100$.
@@
Figure 8: Comparison of approximation based on Kullback-Leibler divergence and total variational distance when the degree of freedom $\nu=2$. The blue line represents the density of $f$. The green line represents the density of approximation $g$ based on total variational distance, and the orange line is the density of approximation $g$ based on Kullback-Leibler divergence.
@@
Let $f(z|\alpha,\nu)$ be the skewed Student $t$ density function, defined as follows:
@@
By definition of Jeffery’s prior, we have $\boldsymbol{\pi}(\alpha|\nu)\propto\sqrt{I_{\alpha\alpha}}$, where $I_{\alpha\alpha}$ is the Fisher information of $\alpha$ in the skewed Student $t$ density function $f(z|\alpha,\nu)$. According to Branco et al. (2013, Proposition 2), the Fisher information of $\alpha$ in $f(z|\alpha,\nu)$ is given by
@@
$H_{6}$ $=\,_{2}F_{1}\left(\frac{1}{2},\nu+2;\frac{\nu+5}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right).$
Furthermore, the detailed proof of $I_{\alpha\alpha}$ is in Lemma A.2 which completes the proof. ∎
Goodness-of-fit for Skewed Student t Distribution
For simplicity, let $x_{1},x_{2},\ldots,x_{n}$ be the observations of $n$ control subjects. We denote $F(x|\alpha,\nu,\xi,\omega)$ as the cumulative function of this distribution where The parameters $\alpha,\nu,\xi,\omega$ are unknown. The definition of skewed Student $t$ distribution is defined in Section 2.2. To test the null hypothesis $H_{0}:x_{1},x_{2},\ldots,x_{n}$ are random samples from the skewed Student $t$ distribution, we propose the following testing process:
@@
More details on the testing procedures and critical points can be found in Chen and Balakrishnan (1995). If we accept the null hypothesis $H_{0}$ at some significance level, we can proceed with comparing a single-subject and the control group described in the next section. If $H_{0}$ is rejected, one can employ a non-parametric method to compare a single-subject with the control group. However, the discussion of the non-parametric method is beyond the scope of this paper. We plan to extend this approach in future work.
@@
Table 7: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under one-sided test (less). Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’, ‘CG-HA’ and ‘AD’ denote Crawford-Garthwaite Bayesian method, Crawford-Garthwaite Bayesian method based on hyperbolic arcsine transformation and Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG-HA CG AD BIGPAST
FPR 50 0.0874 0.0788 0.0492 0.0788 0.2214 0.0473
100 0.0677 0.0630 0.0419 0.0630 0.1920 0.0389
200 0.0595 0.0571 0.0265 0.0571 0.1972 0.0223
400 0.0533 0.0520 0.0329 0.0520 0.1870 0.0144
TPR 50 0.9631 0.9574 0.9367 0.9574 0.8090 0.9190
100 0.9820 0.9802 0.9779 0.9802 0.8393 0.9571
200 0.9974 0.9968 0.9980 0.9968 0.8586 0.9790
400 0.9983 0.9982 0.9898 0.9982 0.8564 0.9819
ACC 50 0.9379 0.9393 0.9437 0.9393 0.7938 0.9359
100 0.9572 0.9586 0.9680 0.9586 0.8237 0.9591
200 0.9689 0.9698 0.9857 0.9698 0.8307 0.9784
400 0.9725 0.9731 0.9784 0.9731 0.8347 0.9837
Table 8: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under a one-sided test(greater). Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’, ‘CG-HA’ and ‘AD’ denote Crawford-Garthwaite Bayesian method, Crawford-Garthwaite Bayesian method based on hyperbolic arcsine transformation and Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG-HA CG AD BIGPAST
FPR 50 0.0003 0.0002 0.0023 0.0002 0.4511 0.0030
100 0.0000 0.0000 0.0020 0.0000 0.4654 0.0044
200 0.0000 0.0000 0.0003 0.0000 0.4699 0.0027
400 0.0000 0.0000 0.0001 0.0000 0.4701 0.0021
TPR 50 0.6402 0.5907 0.9103 0.5908 0.8229 0.8858
100 0.5445 0.5197 0.8975 0.5198 0.7792 0.9206
200 0.5824 0.5693 0.9138 0.5694 0.8491 0.9628
400 0.5389 0.5323 0.9281 0.5323 0.8197 0.9670
ACC 50 0.8199 0.7953 0.9540 0.7953 0.6859 0.9414
100 0.7722 0.7599 0.9478 0.7599 0.6569 0.9581
200 0.7912 0.7847 0.9568 0.7847 0.6896 0.9800
400 0.7694 0.7661 0.9640 0.7661 0.6748 0.9825
Table 9: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 20% positive and 80% negative, are presented under a two-sided test. Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’, ‘CG-HA’ and ‘AD’ denote the Crawford-Garthwaite Bayesian method, the Crawford-Garthwaite Bayesian method based on hyperbolic arcsine transformation and the Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG-HA CG AD BIGPAST
FPR 50 0.1383 0.1237 0.0498 0.1237 0.1170 0.0492
100 0.1341 0.1261 0.0485 0.1262 0.0699 0.0411
200 0.1218 0.1174 0.0367 0.1174 0.0641 0.0303
400 0.1294 0.1273 0.0396 0.1273 0.0468 0.0268
TPR 50 0.6622 0.6131 0.8243 0.6130 0.8519 0.8446
100 0.6439 0.6222 0.9149 0.6222 0.9589 0.9270
200 0.6275 0.6160 0.9165 0.6161 0.9650 0.9327
400 0.6036 0.5978 0.9196 0.5979 0.9805 0.9761
ACC 50 0.8218 0.8236 0.9250 0.8236 0.8768 0.9295
100 0.8215 0.8235 0.9442 0.8235 0.9358 0.9525
200 0.8281 0.8293 0.9539 0.8293 0.9417 0.9623
400 0.8172 0.8177 0.9523 0.8178 0.9586 0.9738
Table 10: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under a two-sided test. The underlying density is the Normal density function. Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
@@
Figure 11: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=1$ and $\nu=10$.  Figure 11(a) shows the densities of skewed Student $t$ and Normal distributions, the $\mu$ and $\sigma^{2}$ of Normal distribution are equal to the mean and variance of skewed Student $t$ distribution. The Normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student $t$ distribution. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.
@@
Figure 12: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=1$ and $\nu=5$.  Figure 12(a) shows the densities of skewed Student $t$ and Normal distributions, the $\mu$ and $\sigma^{2}$ of Normal distribution are equal to the mean and variance of skewed Student $t$ distribution. The Normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student $t$ distribution. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.
@@
Figure 13: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=2$ and $\nu=5$.  Figure 13(a) shows the densities of skewed Student $t$ and Normal distributions, the $\mu$ and $\sigma^{2}$ of Normal distribution are equal to the mean and variance of skewed Student $t$ distribution. The Normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student $t$ distribution. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.